OpenStudy (anonymous):

There are two fruit trees located at (3, 0) and (−3, 0) in the backyard plan. Maurice wants to use these two fruit trees as the focal points for an elliptical flowerbed. Johanna wants to use these two fruit trees as the focal points for some hyperbolic flowerbeds. Create the location of two vertices on the y-axis. Show your work creating the equations for both the horizontal ellipse and the horizontal hyperbola. Include the graph of both equations and the focal points on the same coordinate plane.

OpenStudy (anonymous):

@Hero plz help

OpenStudy (anonymous):

i can help you :)

OpenStudy (anonymous):

OMG THANK YOU SO MUCH I NEED SLEEP OMG

OpenStudy (anonymous):

lol ok :)

OpenStudy (anonymous):

(3,0) and (–3,0) are to be the focal points of the ellipse right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

x^2/25+y^2/16=1 is that the equation for the elliptical flower bed?

OpenStudy (anonymous):

for this you want \[\frac{ x2 ^{} }{ a2 } + \frac{ y2 }{ b2 } = 1\] and you want \[a^{2} - b^{2} = 3^{2}\] the easiest way to do that is to use the famous 3−4−5 right triangle and make a=5,b=4 so c=3 and use \[\frac{ x^{2} }{ 5^{2} } + \frac{ y^{2} }{ 4^{2} } = 1\] that will make your foci (−3,0) and (3,0)

OpenStudy (anonymous):

ok I got that part but what about the hyperbolic flowerbed

OpenStudy (anonymous):

can you try

OpenStudy (anonymous):

I can't make up the value of A

OpenStudy (anonymous):

like I know x^2/a^2-y^2/b^2=1 and a^2+b^2=3^2

OpenStudy (anonymous):

for the hyperbola it will look like x2/ a2−y2/b2=1 and you want a2+b2=33 simplest way i can think to do it is to make a2=8,b2=1 and use x2/8−y2=1 but you have other choices

OpenStudy (anonymous):

how did you get 33?

OpenStudy (anonymous):

no its \[3^{3}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

medal and fan please

OpenStudy (anonymous):

yeah but one more question how do you graph the second one?

OpenStudy (anonymous):

here is a nice graph with both together if you need one http://www.wolframalpha.com/input/?i=+x^2%2F8-y^2%3D1%2Cx^2%2F25%2By^2%2F16%3D1

OpenStudy (anonymous):

ok thank you so much!!

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

medal and fan?

OpenStudy (anonymous):

yeah i already did

OpenStudy (anonymous):

ok thanks :)